• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 22:14:54 on July31,2005

    Entry number 150242

    keyword=Lost many ROCs... aargh...

    
    At 20:45 we lost, all at the same time, communications to four
    IOCs:
    
     We see that the later three are gone from the white screens on
    the HALLA_O_ioc.adl GUI.
    
    Ask MCC for no beam. End HAPPEX run 4004. 
    
    Some sort of network problem? We hear Bob Michael's voice
    telling us that the sweeper magnet current is too low (because
    lost communication with iocha49....) To pay him back for his concern, 
    we page him.   :-)
    
    Target IOC needed to be rebooted. Reboots OK. 
    
    IOCHA49 came back on its own.
    
    IOCHA14 was rebooted through the web via  hareboot10.
    
    No WWW reboot interface for IOCHA16. How to reboot? Discover, after some 
    scrambling the reboot button in rack in the middle room. Reboot.
    The green button is labelled "right spectrometer magnet control VME #16" 
    in 
    Reset Switchbox #1.
    
    After a little bit we find, naturally, that we have lost Q2, Q3 on
    each arm, and that the diples are at the wrong fields. 
    
    Reset, then ON for the quadrupoles. 
    
    Set P0 to get the dipoles ramping. 
     
    Cycle Q2 and Q3 on each arm.
    
    Adjust septa to 10% settings. 
    
    But, we have a hard time getting the dipoles locked...
    
    Finally, after GOOGLE'ing the JLab WWW site (thanks for the 
    idea Bob!), we find a useful log entry from last year from
    Jaideep, halog entry 129538, which gives instructions that eventually work. 
    I will put these into the HAPPEX How-To instructions. Thanks Jaideep!
    
    21:50 - all OK again, ask for 35 µA beam, and start new HAPPEX run 
    (run 4005).